### v2

@@ Abstract.
Security engineering, from security requirements engineering to the implementation of cryptographic protocols, is often supported by domain-specific languages (DSLs). Unfortunately, a lack of knowledge about these DSLs, such as which security aspects are addressed and when, hinders their effective use and further research. This systematic literature review examines 120 security-oriented DSLs based on six research questions concerning security aspects and goals, language-specific characteristics, integration into the software development lifecycle (SDLC), and effectiveness of the DSLs. We observe a high degree of fragmentation, which leads to opportunities for integration. We also need to improve the usability and evaluation of security DSLs.
domain specific languages, security, survey
ccs: Security and privacyccs: Software and its engineering Domain specific languages
@@ 1. Introduction
Security is a fundamental aspect to be considered in the engineering of software-intensive systems, as security breaches can have serious consequences. Consider an Electronic Health Records (EHR) system such as iTrust (Heckman et al., 2018; Meneely et al., 2021): It is important that sensitive patient information is not exposed to unauthorized entities, but it would be even worse if an attacker were able to manipulate medical records, potentially leading to incorrect and harmful treatments. Although there is a relatively high level of awareness among developers today about the security of the software-intensive systems they develop (Xie et al., 2011), successful attacks on such systems are still frequently reported, such as hospitals being forced to shut down due to an attack on their EHR systems (Singh and Sittig, 2016).
@@
One of the main challenges in developing secure software-intensive systems is that security engineering spans the entire Software Development Life Cycle (SDLC) (McGraw, 2004). Security has to be considered with entirely different levels of abstraction throughout the entire SDLC and security compliance among all developed artifacts must be ensured (Peldszus, 2022). However, already properly planning or realizing security on a single artifact is non-trivial. Relevant security requirements are often implicit (Bartsch, 2011), which hardens to address these systematically. Furthermore, common development artifacts, such as source code that is usually written in general purpose programming languages, do not support easy reasoning about security but security experts should be given system representations that are tailored towards their tasks (Peldszus, 2022).
Contrary to general purpose programming languages, Domain-Specific Languages (DSLs) (Wasowski and Berger, 2023) are purpose-built to cater to the needs of a particular problem domain. While DSLs can encompass programming languages, they can also target other development phases, such as design-time system descriptions. Leveraging the inherent specificity of the target domain, DSLs empower developers to address problems and engineer solutions at the most appropriate level of abstraction. By providing a more intuitive representation of systems, DSLs simplify the design process. Moreover, DSLs often offer the capability to expose parameters that are challenging to express with general-purpose languages. Furthermore, DSLs are crucial for leveraging powerful supporting tools, enabling optimal design exploration, automated testing, and formal verification of specific properties.
@@
Due to their power, domain specific languages have been proposed for facilitating a wide range of security engineering tasks. In general, information security refers to the practice of protecting information from unauthorized access, disclosure, modification, or destruction. It involves a set of measures and techniques designed to ensure the three core principles: confidentiality, integrity, and availability of information (Cawthra et al., 2020). Multiple goals can be derived from these core principles or are strongly related, including authentication, privacy and accountability. To counter the many possible attack vectors, which threaten security goals, a wide range of defensive techniques and building blocks exist at various system levels. For example, access policies form a fundamental component, establishing guidelines and controls to determine who is granted access to specific resources within an organization’s network. At a lower level, cryptography provides an existential toolkit to enforce security policies by preventing malicious parties from reading and manipulating sensitive data.
DSLs play a critical role in addressing the unique requirements of specific domains—including security in particular. The scientific community has contributed a variety of security DSLs, targeting different phases of the SDLC and different aspects of security. This broad security spectrum is reflected in the highly fragmented research on security DSLs, which limits the widespread adoption of security DSLs and hinders their development. This survey aims to provide a comprehensive overview of DSLs in the scientific literature, focusing on those with security-specific language features. Our goal is to identify connections and gaps, and to suggest possible combinations and integrations. In particular, there is currently a lack of work that provides a holistic analysis of DSLs across different security domains. We make a first attempt to organize previous research in the area of DSLs for security. The following research questions guide our investigation:
@@ 1. RQ1:
    What Security DSLs have been presented in the scientific literature?
    While the scientific community has contributed numerous security DSLs, there is currently no systematic overview of them. We will explore the existing literature and focus on DSLs that address security concerns in software development to provide such a missing catalog of security DSLs.
@@ 2. RQ2:
    What security aspects are targeted by the security DSLs?
    As discussed, security DSLs have been designed to address different aspects of security in terms of different security goals, attacker models, and defense mechanisms. In doing so, they can be problem-oriented, focusing on general security planning and the identification of potential threats, or solution-oriented, facilitating the realization of concrete defense mechanisms. To this end, we aim to capture these aspects in order to contextualize the security DSLs.
@@ 3. RQ3:
    What phases of the development process are supported by the DSLs?
    In the context of developing secure software-intensive systems, security DSLs need to be integrated into the development process. Thereby, DSLs can be relevant for a single development phase, but also for several phases. Accordingly, we aim to identify which SDLC phases are relevant for each security DSL and which security aspects the DSLs focus on in which phases of the software development lifecycle. In particular, for DSLs that aim to support multiple phases, we are interested in which phase they are instantiated and in which phases they are used.
@@ 4. RQ4:
    What types of DSLs exist, and how are they used?
    When designing a DSL, language engineers can choose between different types of DSLs with different characteristics. To provide an overview of the existing DSLs that can serve as a reference, we will also examine their general characteristics, such as whether they are stand-alone (external) DSLs or embedded in other established languages (internal DSL), such as a general-purpose programming language, and whether they are textual or visual DSLs. In addition, we will examine the tools associated with these DSLs to understand how they are used in practice.
@@ 5. RQ5:
    What are the semantics of the DSLs?
    Based on the targeted security goals and the SDLC phases in which the DSLs are addressing these goals, we investigate how this is semantically realized. Specifically, we are interested in what mechanism brings the DSLs instances into effect. To this end, we aim at capturing how are the instances of the DSLs executed to achieve the corresponding security goals or what kind of transformation is provided by the backend to propagate and utilize the security properties specified using the DSLs. We will also explore how the DSLs are integrated into the development process.
@@ 6. RQ6:
    How effective are the DSLs?
    Lastly, we are interested in how effective the DSLs are and how their effectiveness has been proofed or validated. Primary, this questions aims at (formal) security guarantees including properties like soundness, precision, and completeness. However, we also are interested in empirical data on the security DSLs and whether the DSLs have been evaluated on real-world user studies.
With this paper we hope to support multiple groups of readers. First, we want to give developers of secure systems an overview of DSLs that can support their work at different stages of the software development life cycle. Our work is also aimed at security researchers who want to leverage DSLs for their specific security domain and are looking for examples, best practices and universal solutions from existing security DSLs. Lastly we target security researchers working on language-based security, possibly across multiple security domains and attacker models, interested in integration, interaction and combination of DSLs.
@@ 2. Background
We first give a general introduction to domain-specific languages, and then discuss information security in the context of domains that need to be considered when developing secure systems.
@@ 2.1. Domain-Specific Languages
A DSL is a programming or specification language that is dedicated to a particular problem domain, problem representation, or solution technique (Mernik et al., 2005; Fowler, 2011; Wasowski and Berger, 2023). DSLs are designed to simplify the development process within a specific context or domain by providing abstractions and syntax that directly represent concepts and operations relevant to that domain. Usually, they hide unnecessary details and complexity present in general-purpose languages, allowing developers to focus on solving problems within their domain of expertise, such as secure software design. Because DSLs are tailored to the needs of a specific domain, they are more expressive, concise, and easier to use within that domain than general-purpose programming languages. To this end, DSLs can greatly improve productivity and code maintainability within specific domains by allowing developers to express solutions using terminology and abstractions familiar to domain experts. However, designing and implementing DSLs requires careful consideration of the domain requirements, the target audience, and the tradeoffs between expressiveness and ease of use.
@@ 2.2. Information Security
Information security aims to control the flow of sensitive information at various system levels (Anderson, 2008). Thereby, the three core objectives of information security are confidentiality, integrity and availability (Cawthra et al., 2020). Confidentiality guarantees that secret information is only accessible to intended parties. Integrity refers to the dependability of information and prevents undesired modification of trusted data. Finally, availability ensures that information and information systems are accessible and operational, preventing disruptions that could result in the unavailability of critical resources. Without the goal of availability, security could be simply achieved by preventing access to sensitive information for any entity. Other security objectives can be derived from the three core objectives: Privacy ensures the confidentiality of private data (Spiekermann, 2012), authenticity ensures the integrity of a message’s content and origin, Accountability ensures the availability and integrity of an operation’s actor and non-repudiability ensures the accountability of a message’s sender or receive (International Organization for Standardization and International Electrotechnical Commission (IEC)(2022), ISO).
2.2.1. Requirements & Threat Modeling
@@
Before analyzing, enhancing, or building a system, the security requirements are determined and inferred from higher-level goals (Türpe, 2017). For example, laws might demand privacy measures that necessitate controls at the enterprise level which again lead to requirements at the software level (Deng et al., 2011).
@@
Requirements can also be identified after examining potential threats to security. Threat modeling (Shostack, 2008) is a fundamental methodology to comprehensively evaluate the security of a system by envisioning potential attack vectors and identifying gaps in the defense. Threat modeling paired with the computation of probabilities enables the assessment of security risks. Attack patterns can also serve for detection during runtime.
After analyzing and planning security measures for a complex system, usually multiple defense mechanism are required to address all potential threats, as the options for an attacker are manifold and can target different system layers and components. A well-defined countermeasure is therefore often accompanied by an attacker model (Rocchetto and Tippenhauer, 2016) that defines the capabilities and limitations of potential attackers it aims to defend against.
2.2.2. Access & Information Flow Control
A fundamental concept to achieve basic information security goals is access control (Sandhu and Samarati, 1994), which limits the access to information for certain entities. The formal foundations can be traced back to the Bell-LaPadula model (Bell and LaPadula, 1973) (confidentiality) and the Biba model (Biba, 1977) (integrity). Access control systems regulate the physical or digital access to places and resources with rules expressed in policies. The access policies can be based on identities, roles, or attributes. Multi-user operating systems, for example, restrict the access to data for users and differentiate between permissions for read, write, and execute.
@@ 2.2.2. Access & Information Flow Control
Similar to access control, but acting at the level of program variables is information flow control (Sabelfeld and Myers, 2003). In particular with indirect dependencies, the flow of sensitive information is difficult to track and protect in complex software systems and errors can lead to violations of security policies. Beginning with the seminal work of Denning and Denning (Denning and Denning, 1977), information flow control tracks the propagation of information in a program and verifies its compliance with policies. These policies ensure either integrity by prohibiting untrusted input to influence trusted output, or confidentiality by enforcing the public output to be independent of secure input.
2.2.3. Cryptography
@@
Eventually, secure systems have to transmit sensitive information over insecure channels, e.g., the internet or store it on unprotected devices. Cryptography with primitives like symmetric or asymmetric ciphers or hash functions allows maintaining security in these circumstances. For example, encryption ensures the confidentiality of private data by making it indistinguishable from random data. With digital signatures (Cheon et al., 2023), the integrity of messages can be verified. A threat to cryptographic implementations are side-channel attacks which use unintended information channels of software and hardware systems to extract sensitive information (Krausz et al., 2023). Well known examples are attacks utilizing timing side-channels (Kocher, 1996) or power side-channels (Kocher et al., 1999). When attacking an implementation using a side-channel, the adversary does not attempt to break the cryptographic function itself but exploits the weaknesses of its implementation and the unintended leakage of information produced by it. For example, differences in execution time of a cryptographic function that are secret dependent, can be exploited to infer the secret itself.
Advanced cryptography includes concepts like secure Multi-Party Computation (MPC) (Goldreich, 1998) and Fully Homomorphic Encryption (FHE) (Gentry, 2009). MPC enables a group of mutually distrustful parties to compute an arbitrary, joint function of their private data. The data is contributed by each party without revealing it to any other party, not even to a trusted third party. FHE even allows arbitrary computations on encrypted data without decryption, thus preserving confidentiality.
2.2.4. Intrusion Detection
@@
Besides preventive measures, security can be enhanced with detection methods. An intrusion detection system tries to identify attacks at runtime. The most prominent examples are network intrusion detection systems observing network packages and host-based intrusion detection systems, which observe data files. Intrusion detection can be classified (Axelsson, 2000) into anomaly and signature (pattern) based approaches.
@@ 3. Methodology
We used a systematic review methodology (Ralph, 2021) to investigate the state of the art in security DSLs. In what follows, we first describe our DSL selection criteria and the corresponding selection process, followed by a description of the analysis methodology of the selected DSLs.
@@ 3.1. DSL Selection
For our systematic review, we searched for security DSLs that can be used by developers, security professionals, and other stakeholders in the development of secure software-intensive systems. We focused on DSLs that explicitly address the security of the system. To this end, we considered all security DSLs that met the following two inclusion criteria
@@
The DSL will be used in the development of a secure software-intensive system.
@@
To systematically identify relevant security DSLs for in-depth analysis, we searched the academic literature on security DSLs. In order to obtain a broad initial coverage of the landscape of security DSLs, we decided to use a generic search term in the initial search and then filter out irrelevant results. Based on the initial dataset of publications identified in this way, we performed an iteration of backward snowballing (Wohlin, 2014) to ensure that we did not miss relevant publications on security DSLs. Figure 1 visualizes the individual sampling and filtering steps performed. In the first step, the initial search, we used the following search term on Semantic Scholar:
security domain-specific language
To avoid dated DSLs that had only low-impact, assure the quality of the papers, and avoid papers from other domains or not focusing on usable security DSLs as aimed by us according to the inclusion criteria, we defined six exclusion criteria.
E1 – published before 2010.:
In our initial search, we excluded papers published before the year 2010. However, we did not apply this exclusion criterion in the snowballing, therefore still including relevant older publications in our data set.
E2 – not in English.:
We excluded papers not written in English for practical reasons.
E3 – low quality.:
To ensure a certain level of scientific quality, we excluded papers that were not peer-reviewed or published in a predatory venue or by a predatory publisher. Please note that the security DSLs themselves must not be the product of scientific papers but must have been discussed in at least one scientific paper.
E4 – not computer science.:
Since we focus on security DSLs in software development, we excluded papers that do not belong to the field of computer science.
E5 – no security DSL.:
We excluded all papers that are not explicitly about security DSLs but only mention these on the side, focus only on languages or only on security but not DSLs.
E6 – no usable DSL.:
Since we aim at practically applicable security DSLs, we excluded papers that do not describe a usable DSL but only theoretical concepts or models.
We searched Semantic Scholar with a python script using the public API of Semantic Scholar and received 14,255 documents. According to exclusion criteria E1, E2 and E4, we limited this search to papers published in 2010 or later, papers in English, and the field of computer science. Based on this accumulation we proceeded with automated filtering, at first removing duplicates, which reduced the number of papers to 811. From an initial manual inspection, we gathered a set of common keywords that indicate research irrelevant for this survey according to exclusion criterion E5:
@@
We then automatically scanned the title and abstract and paper for these keywords and reduced the papers to 670. From these results we continued with a manual selection based on the title and abstract. We omitted remaining papers not written in English (E2). Furthermore, we looked for indications that the paper either presents a security DSL, discusses the usability or integration of an existing DSLs or surveys DSLs in a security domain and excluded all papers not meeting on of these three (E5 and E6). With 133 papers left we narrowed down the set with a manual selection based on the full paper. In doing so, we also discarded papers, that were not published at a scientific conference or journal, in particular, we discarded three papers, which were solely published as preprints. We rejected three papers, which were published at predatory journals according to Beall’s list [^1] Beall’s list of predatory publishers: https://beallslist.net/ (E3). Multiple publications were excluded, simply because they do not cover a security DSL (E5), although the title and abstract gave indications for this. We excluded several papers, that presented type systems or formal calculi without a usable language (E6). We also dismissed languages and tools that, although used in a security context, come without security specific language features, but for example only a backend which provides security features. Prepose (Figueiredo et al., 2016) is an example for this, as the system has clearly a security goal, but the DSL is completely security agnostic. For three papers, we were not able to access the full document and therefore omitted them.

### v3

@@ Abstract.
Security engineering—from creating security requirements to the implementation of security features, such as cryptography or authentification—is often supported by domain-specific languages (DSLs). While many security DSLs have been presented, a lack of overview and empirical data about these DSLs, such as which security aspects are addressed and when, hinders their effective use and further research. This systematic literature review examines 120 security DSLs regarding their security aspects and goals addressed, their language-specific characteristics, their integration into the software development lifecycle, and their evaluation. We observe a focus on individual development phases and a high degree of fragmentation, which leads to opportunities for integration. The research community also needs to improve the usability and evaluation of security DSLs.
copyright: noneccs: Security and privacyccs: Software and its engineering Domain specific languages
@@ 1. Introduction
Information security is a fundamental aspect when engineering software systems, as security breaches can have serious consequences. Consider an Electronic Health Records (EHR) system, such as iTrust (Heckman et al., 2018; Meneely et al., 2021). It is crucial that sensitive patient information is not exposed to unauthorized entities, but it would be even worse if an attacker could manipulate medical records, potentially leading to incorrect and harmful treatments. Although there is a relatively high level of awareness among developers about the security of software they develop (Xie et al., 2011; Hermann et al., 2025a), successful attacks are still frequently reported, such as hospitals being forced to shut down due to attacks on their EHR systems (Singh and Sittig, 2016).
@@
A core challenge is that security spans the entire software development life cycle (SDLC) (McGraw, 2004). Security has to be considered at different levels of abstraction while ensuring compliance among all developed artifacts (Peldszus, 2022). However, already planning or realizing security on a single artifact is non-trivial and challenges developers (Hermann et al., 2025a). Relevant security requirements are often implicit (Bartsch, 2011), which makes addressing them hard. Furthermore, common artifacts, such as source code, do not support easy reasoning about security. Security experts and other engineers should be provided with system representations that are tailored towards their tasks (Peldszus, 2022).
Contrary to programming languages, which are general-purpose, domain-specific languages (DSLs) (Wasowski and Berger, 2023) are purpose-built and address the needs of a particular problem domain. DSLs can support the implementation of security features (Peldszus, 2022; Hermann et al., 2025a) as well as other phases, such as designing software systems. While DSLs are less expressive than general-purpose languages, they support developers and other engineers (e.g., security experts) by providing appropriate abstractions expressed in domain-specific terminology in one language (Wasowski and Berger, 2023). For instance, DSLs can represent software at a higher level of abstraction, removing irrelevant details, making them more intuitive and easier to understand. They can simplify the design or reasoning about specific aspects, such as security.
@@
This survey focuses on DSLs designed for information security, which aims at protecting information from unauthorized access, disclosure, modification, or destruction. To counter the many possible attack vectors, a wide range of defensive techniques and building blocks exist at various system levels. For example, access policies form a fundamental component, establishing guidelines and controls to determine who is granted access to specific resources. At a lower level, cryptography enforces security policies by preventing malicious parties from reading and manipulating sensitive data.
The scientific community has contributed numerous security DSLs, which focus on specific security aspects and that offer security-specific language constructs. Such DSLs often target a particular phase of the SDLC. As we will show, the DSLs cover a broad spectrum of security, while research is highly fragmented. The lack of an overview and systematic empirical data on security DSLs limits the widespread adoption of security DSLs, hinders their development, and prevents future research.
This survey provides a comprehensive overview of security DSLs discussed in peer-reviewed academic literature. Our goal is to identify patterns, trends, and open problems, and to suggest possible combinations of complementary security DSLs and integrations into the SDLC. We organize the existing body of research on security DSLs, guided by the following research questions.
@@ 1. RQ1:
    What security DSLs have been presented in the scientific literature? While researchers contributed many security DSLs, no systematic overview exists. We explore the existing literature and focus on DSLs that address security concerns in software development.
@@ 2. RQ2:
    What security aspects are targeted by security DSLs? Security DSLs were designed to address different aspects of security, such as security goals, attacker models, and defense mechanisms. As such, they can be problem-oriented, focusing on general security planning and the identification of potential threats, or solution-oriented, facilitating the realization of defense mechanisms. We capture these security aspects to contextualize the security DSLs.
@@ 3. RQ3:
    What phases of the development process are supported by security DSLs? When developing secure software systems, security DSLs need to be integrated into the development process. DSLs can be relevant for a single or multiple development phases. We identify which SDLC phases are relevant for each security DSL, whether there are any phases particularly addressed by security DSLs, and which security aspects the DSLs focus on in which phases. Specifically for DSLs that support multiple phases, we determine in which phase they are instantiated and in which they are used.
@@ 4. RQ4:
    What types of security DSLs exist, and how are they used? When designing a DSL, language engineers can choose between different types of DSLs, each type with its own characteristics. We examine these characteristics, such as whether a DSL is stand-alone (external) or embedded in an established language, such as a programming language, (internal) and whether a DSL is textual or visual. We also examine the tools associated with these DSLs to understand how they are used in practice.
@@ 5. RQ5:
    What are the semantics of security DSLs? Based on the targeted security goals and the targeted SDLC phases, we investigate how these are semantically realized, specifically what mechanism bring the DSL instances into effect. We capture how instances of security DSLs are executed or interpreted to achieve their security goals, or what kind of transformation is provided by the backend to propagate and utilize any security properties specified using the DSL. We also explore how security DSLs are integrated into the development process.
@@ 6. RQ6:
    How are security DSLs evaluated? Finally, we determine whether and how the DSLs have been evaluated, particularly their effectiveness. We analyze what security guarantees including properties like soundness, precision, and completeness were considered and perhaps even proven. We also identify and report quantitative and qualitative empirical data provided for the DSLs.
Our study supports developers, researchers, and tool vendors. Developers and researchers obtain an overview of DSLs to support their work at different SDLC stages. They learn about examples, best practices, and universal solutions from existing security DSLs. Researchers and tool vendors, such as those working on language-based security, possibly across multiple security domains and attacker models, learn about security DSL concepts and about their integration, interaction, and combination.
@@ 2. Background
We introduce the necessary background on DSLs and information security, together with examples.
@@ 2.1. Domain-Specific Languages
A DSL is a programming, modeling, or specification language dedicated to a particular problem domain, problem representation, or solution (Mernik et al., 2005; Fowler, 2011; Wasowski and Berger, 2023). DSLs simplify the development process within a specific context or domain by providing abstractions and syntax that directly represent concepts and operations relevant to that domain. DSLs typically hide unnecessary details and complexity present in general-purpose languages, allowing developers to focus on solving problems within their domain of expertise, such as software security. Since DSLs are tailored to the needs of a domain, they are more intuitive, concise, and easier to use than general-purpose languages. To this end, DSLs can improve productivity and code maintainability within specific domains by allowing developers to express solutions using terminology and abstractions familiar to domain experts. However, designing and implementing DSLs requires careful consideration of the domain and the target users, making tradeoffs between expressiveness and ease of use.
@@ 2.2. Information Security
Information security controls the flow of sensitive information at various system levels (Anderson, 2008) with three core objectives: confidentiality, integrity, and availability (Cawthra et al., 2020). Confidentiality guarantees that information is only accessible to intended parties. Integrity prevents undesired modification of trusted data. Availability prevents disruptions that could result in critical resources being unavailable. Furthermore, privacy ensures confidentiality of private data (Spiekermann, 2012); authenticity ensures the integrity of messages’ content and origin; accountability ensures availability and integrity of actors; and non-repudiability ensures the accountability of messages’ sender or receiver (International Organization for Standardization (ISO) and International Electrotechnical Commission (IEC), 2022).
@@
Before analyzing, enhancing, or building a system, the security requirements are usually determined and inferred from higher-level goals (Türpe, 2017). For example, laws might demand privacy measures that necessitate controls at the enterprise level, which imply software requirements (Deng et al., 2011).
@@
Requirements can also be identified after examining potential security threats. Threat modeling(Shostack, 2008) is a method to comprehensively evaluate the security of a system by envisioning attack vectors and identifying gaps in the defense. Threat modeling combined with the computation of probabilities enables proper assessment of security risks. Attack patterns can also serve for detection during runtime. After analyzing and planning, usually multiple defense mechanism are required to address the threats, since attackers have many options and can target different system layers and components. So, countermeasure are often accompanied by an attacker model(Rocchetto and Tippenhauer, 2016) that defines the capabilities and limitations of potential attackers to defend against.
@@
Access control(Sandhu and Samarati, 1994) limits access to information, with foundations going back to the Bell-LaPadula (Bell and LaPadula, 1973) (confidentiality) and the Biba model (Biba, 1977) (integrity). Access control regulates physical or digital access to resources expressed as policies, which can be based on identities, roles, or attributes. Multi-user operating systems, for example, manage read, write, and execute permissions for users. Information flow control(Sabelfeld and Myers, 2003) is similar to access control, but at the level of program variables. Information flow is often hard to track and protect in complex systems, and errors can lead to policy violations. Flow control tracks data propagation in programs and verifies compliance with policies (Denning and Denning, 1977), which ensure integrity by prohibiting untrusted input to influence trusted output, or confidentiality by enforcing public output to be independent of secure input.
@@
When transmitting sensitive data over insecure channels or storing it on unprotected devices, cryptography with primitives (e.g., symmetric and asymmetric ciphers, or hash functions) maintains security. With digital signatures (Cheon et al., 2023), the integrity of messages can be verified. A threat, however, are side-channel attacks, which use unintended information channels to extract sensitive information (Krausz et al., 2023). Examples are timing (Kocher, 1996) and power side-channel (Kocher et al., 1999) attacks. In such attacks, adversaries do not attempt to break the cryptographic function, but exploit implementation weaknesses and unintended information leakage. For example, differences in execution time of a cryptographic function can be exploited to infer secrets. Advanced cryptography includes concepts such as secure Multi-Party Computation (MPC)(Goldreich, 1998) and Fully Homomorphic Encryption (FHE)(Gentry, 2009). MPC enables mutually distrustful parties to compute an arbitrary joint function of their private data, which is contributed by each party without revealing it to others. FHE even allows arbitrary computations on encrypted data without decryption, preserving confidentiality.
@@
Besides preventive measures, security can be enhanced with intrusion detection methods to identify attacks at runtime. They classify (Axelsson, 2000) into anomaly and signature (pattern)-based approaches. Prominent examples are network and host-based intrusion detection.
@@ 3. Methodology
Following ACM SIGSOFT’s Empirical Standards (Ralph, 2021), we used a systematic review methodology to study security DSLs. In what follows, we describe our DSL selection criteria and the corresponding selection process, followed by the analysis methodology of the selected DSLs.
@@ 3.1. DSL Selection
We searched for security DSLs that can be used by developers, security professionals, and other stakeholders for developing secure software systems. We focused on DSLs that were explicitly designed for security, specifically those that meet the following inclusion criteria
@@
The DSL can be used to address information security in software system engineering.
@@
To systematically identify relevant DSLs from the academic literature (cf. I3) and to obtain a broad coverage of the landscape, we decided to use a generic search term in the initial search and then filter out irrelevant results. Based on the initial set of publications identified, as recommended by ACM SIGSOFT’s Empirical Standards (Ralph, 2021), we performed an iteration of backward snowballing (Wohlin, 2014) to ensure that we did not miss relevant publications. Figure 1 visualizes our sampling and filtering steps. The first step (initial search) used the following search term on Semantic Scholar: “security domain-specific language”. To avoid dated DSLs with low impact, to ensure the quality of the papers, and to prevent papers outside our scope, we defined six exclusion criteria.
E1.:
We excluded papers published before 2010 in our initial search. However, we did not apply this criterion for snowballing, so relevant older papers are still in our dataset.
E2.:
We excluded papers not written in English.
E3.:
We excluded non-peer-reviewed papers as well as papers published by a predatory venue or publisher. Note that the security DSLs themselves do not have to be the contribution of scientific papers, but must have been discussed in at least one.
E4.:
We excluded papers that do not belong to the field of computer science.
E5.:
We excluded papers that do not focus on security DSLs and only mention them on the side.
E6.:
We excluded papers without a fully implemented DSL, but only theoretical concepts or models.
We searched Semantic Scholar using its public API and obtained 14,255 documents. According to the exclusion criteria E1, E2, and E4, we limited the search to papers published 2010 or later, written in English, and the field of computer science. Based on this set, we proceeded with automated filtering, first removing duplicates, which reduced the number of papers to 811. From an initial manual inspection, we collected common keywords that indicate research irrelevant to us according to E5:
@@
We then automatically scanned title, abstract, and paper for these keywords and reduced the papers to 670. From these results, we continued with a manual selection based on the title and abstract. We omitted remaining papers not written in English (E2). Furthermore, we looked for evidence that the paper either presents a security DSL, discusses the usability or integration of an existing DSL, or surveys DSLs in a security domain, and excluded papers not meeting any of these three criteria (E5 and E6). With 133 papers left, we narrowed down the set with a manual selection based on the full paper. In doing so, we also discarded papers that were not published at a scientific conference or journal (e.g., three papers solely published as preprints). We rejected three papers published at predatory journals according to Beall’s list (Beall, 2024) (E3). Multiple publications were excluded simply because they did not cover a security DSL (E5), despite title and abstract indicating it. We excluded several papers that presented type systems or formal calculi without a usable language (E6). We also dismissed languages and tools that, although used in a security context, came without security-specific language features, but for example only a backend that provides security features. Prepose(Figueiredo et al., 2016) is such an example, as the system clearly has a security goal, but the DSL is security agnostic. For three papers, we could not access the full document and therefore omitted them.